OpenStudy (anonymous):

Prove by mathematical induction x^(2n)-y^(2n) is divisible by (x+y) for integers n>=1.

OpenStudy (anonymous):

damn wtf. are you doing hsc 2014?

OpenStudy (helder_edwin):

u have to prove that \[\large (x+y)\mid(x^{2n}-y^{2n}) \] for all integers \(n\geq1\)

OpenStudy (anonymous):

yeah bro, 2 days to final, u up for that?

OpenStudy (anonymous):

how helper_irwin kind sir?

OpenStudy (helder_edwin):

(i) induction base: n=1 \[\large x^{2n}-y^{2n}=x^2-y^2=(x+y)(x-y) \] which shows that \((x+y)\mid(x^2-y^2)\).

OpenStudy (anonymous):

wait its that easy?, thought there's a need to prove n=k+1 something like that

OpenStudy (anonymous):

no there is more steps, hes just typing give him time

OpenStudy (anonymous):

thanks man

OpenStudy (loser66):

@SydneySider wait for him, next step will be it

OpenStudy (anonymous):

thank you loser66 heyimtrash!

OpenStudy (helder_edwin):

(ii) inductive step: let \(n>1\) and suppose that (inductive hypothesis) \[\large (x+y)\mid(x^{2n}-y^{2n})\quad\text{i.e.}\quad x^{2n}-y^{2n}=(x+y)\cdot\alpha. \] we have to prove that (inductive thesis) \[\large (x+y)\mid(x^{2(n+1)}-y^{2(n+1)})\quad\text{i.e.}\quad x^{2(n+1)}-y^{2(n+1)}=(x+y)\cdot\beta. \]

OpenStudy (anonymous):

thanks bro

OpenStudy (anonymous):

what class u in btw?

OpenStudy (helder_edwin):

let's see: from the inductive hypothesis: \[\large x^{2n}=y^{2n}+(x+y)\alpha \] so \[\large x^{2(n+1)}-y^{2(n+1)}=x^{2n}x^2-y^{2n}y^2 \] \[\large =[y^{2n}+(x+y)\alpha]x^2-y^{2n}y^2 \] \[\large =y^{2n}x^2+(x+y)\alpha x^2-y^{2n}y^2 \] \[\large =y^{2n}(x^2-y^2)+(x+y)\alpha x^2 \] \[\large =y^{2n}(x+y)(x-y)+(x+y)\alpha x^2 \] \[\large =(x+y)[y^{2n}(x-y)+\alpha x^2] \] which proves that \((x+y)\mid)(x^{2(n+1)}-y^{2(n+1)})\)

OpenStudy (anonymous):

thank you so much good man

OpenStudy (loser66):

@helder_edwin thanks a lot. It's quite new to me ( I mean the way to prove) but it's perfect

OpenStudy (anonymous):

i got stuck at the x^2nx^2 part :( damn. but very nice. thanks alot